# On the Equivalence of Random Network Distillation, Deep Ensembles, and Bayesian Inference

### v1

@@ B.2 Posterior Equivalence
B.2.1 Proof of Proposition 4
[…]
@@ B.2.1 Proof of Proposition 4
The proof will show that in the limit $n\to\infty$ the function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ converges to a GP $\tilde{g}(x;\vartheta_{0},\psi_{0})\sim\mathcal{GP}(0,\Theta^{\leq L-1}_{xx^{\prime}})$ by Lévy’s continuity theorem, which we recall informally below. {restatable}thmlevycontinuity(Lévy’s continuity theorem) Let $\{Z_{n}\}_{n=1}^{\infty}$ be a sequence of $\mathbb{R}^{n}$-valued random variables. Their characteristic functions $\varphi_{Z_{n}}(t)$ for some $t\in\mathbb{R}^{n}$ are given by
@@
whose characteristic function is $\varphi_{Z}(t)=\varphi(t)$ Rigorous proof can be found for example in durrett2019probability.
@@
B.2.2 Proof of Theorem 4
[…]

### v2

@@ B.2 Posterior Equivalence
B.2.1 Proof of Proposition 4.1
We restate Proposition 4.1 for convenience. See 4.1
@@
The proof will show that in the limit $n\to\infty$ the function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ converges to a GP $\tilde{g}(x;\vartheta_{0},\psi_{0})\sim\mathcal{GP}(0,\Theta^{\leq L-1}_{xx^{\prime}})$ by Lévy’s continuity theorem, which we recall informally below.
Theorem B.4.
(Lévy’s continuity theorem) Let $\{Z_{n}\}_{n=1}^{\infty}$ be a sequence of $\mathbb{R}^{n}$-valued random variables. Their characteristic functions $\varphi_{Z_{n}}(t)$ for some $t\in\mathbb{R}^{n}$ are given by
@@
whose characteristic function is $\varphi_{Z}(t)=\varphi(t)$
Rigorous proof can be found for example in durrett2019probability.
@@
B.2.2 Proof of Theorem 4.2
We restate Proposition 4.1 for convenience. See 4.2
@@
Appendix C Additional Experimental Details
We report additional experimental details and evaluations. As outlined in the main text, we use two-layer fully connected neural networks with SiLU activations and NTK parametrization. All weights and biases are initialized as $\theta\sim\mathcal{N}(0,I)$. We use an ensemble of 512 models and a single multiheaded RND network with 512 heads. A synthetic dataset is generated with $N=10$ train and $\tilde{N}=5000$ test samples from an isotropic Gaussian $x\sim\mathcal{N}(0,I_{3})$. We label training samples with a synthetic target function
$y(x)=x^{0}+x^{1}+x^{2}-2\prod_{i=1}^{3}x^{i},$ (89)
where $x^{i}$ denotes the $i$-th component of vector $x$. All models are trained according to the algorithms outlined in the main text. For this, we use full-batch gradient descent with a learning rate of $0.1$ for all models. Fig. 2 shows additional results of the same experiment, in which we plot individual test-set ensemble variances against RND errors. As the network width increases, ensemble variances and self-predictive RND errors become more correlated and well-calibrated in scale.
Code for full reproduction will be released upon publication.
(a)                          (b)
Figure 2: (a) Scatter plot of test-set errors between predictive variances of ensembles and self-predictive errors of RND. As width increases, errors become more correlated and correctly calibrated in scale. (b) Likewise, for Bayesian ensembles and Bayesian RND.